Tuoo. White dots are 1 millimeter apart on a black paper they are viewed by eye of Pupil diameter 3 millimetre approximately what is the maximum distance at which these dots can be resolved by an eye

Grade:12th pass

1 Answers

Arun
6 years ago
y/D >= 1.22λ/d
Dλ = 10–3 * 3 * 10–3/(1.22* 5 * 10–7)
= 30/6.1 = 5 m
Dmax = 5 m
Regards
Arun (askIITians forum expert)